# When Indemnity Insurance Fails: Parametric Coverage under Binding Budget and Risk Constraints

### v3

@@ Abstract
In high-risk environments, traditional indemnity insurance is often unaffordable or ineffective, despite its well-known optimality under expected utility. We compare excess-of-loss indemnity insurance with parametric insurance within a common mean-variance framework, allowing for fixed costs, heterogeneous premium loadings, and binding budget constraints. We show that, once these realistic frictions are introduced, parametric insurance can yield higher welfare for risk-averse individuals, even under the same utility objective. The welfare advantage arises precisely when indemnity insurance becomes impractical, and disappears once both contracts are unconstrained. Our results help reconcile classical insurance theory with the growing use of parametric risk transfer in high-risk settings.
@@ 1 Introduction
Insurance markets are increasingly strained in regions exposed to low-frequency, high-severity risks such as floods, wildfires, and cyclones. In many such areas, indemnity insurance has become prohibitively expensive or altogether unavailable, as rising hazard intensity, capital requirements, and fixed costs push premiums and deductibles to levels that leave households effectively uninsured. A key practical limitation is that household indemnity insurance is typically written on a full-value (sum-insured) basis: the premium is charged to cover the entire exposure, and the main lever to restore affordability is to raise the deductible. In high-risk areas, this mechanism can push deductibles to economically unrealistic levels (e.g., tens or hundreds of thousands of dollars), rendering the contract formally available but delivering negligible effective protection. By contrast, parametric insurance is inherently scalable: households can purchase a smaller, explicitly bounded layer of protection (a fixed payment per event) without having to insure the full loss. Recent developments in flood and wildfire insurance markets illustrate this trend: insurers have withdrawn from parts of the California residential market, flood insurance take-up remains persistently low in high-risk zones, and coverage gaps are expanding even in advanced economies (e.g., Kousky and Cooke, 2012). Yet when disasters occur, losses are rarely borne privately alone. Instead, they are often socialized ex post through government relief, reconstruction grants, and implicit guarantees. This growing disconnect between private insurance provision and public risk-bearing raises a fundamental economic question: when full indemnification is no longer feasible, what forms of risk transfer can still improve individual welfare?
Classical insurance theory provides a clear benchmark. Under broad and well-understood conditions, an excess-of-loss indemnity contract maximizes the expected utility of a risk-averse agent when premiums are loaded proportionally to expected losses (Arrow, 1971; Raviv, 1979). This result has shaped both theoretical and practical views of optimal insurance design for decades. However, it relies on assumptions that are increasingly violated in high-risk environments. In particular, it abstracts from differences in risk borne and fixed costs borne by insurers across contract types, as well as binding budget constraints faced by households. When insurance is expensive, the deductible that is still affordable becomes so large that indemnity insurance offers little or no effective protection, even though it remains optimal in a purely formal sense.
@@
Nevertheless, indemnity insurance is still widely considered the best (or only) way to provide effective protection. We show that this view is incomplete. Using a tractable model of frequency and severity, we compare indemnity and parametric insurance under the same expected mean–variance utility criterion. Losses arrive according to a Poisson process and have censored exponential severity. This specification can capture the salient features of natural disaster risk (low-frequency, high-severity) while allowing for closed-form solutions. The indemnity contract is of excess-of-loss type with a deductible, while the parametric contract pays a fixed amount per event and shares the same trigger, so there is no basis risk on frequency. Importantly, we allow for distinct premium loadings and fixed costs across contract types, reflecting the higher capital requirements, claims management costs, and reconstruction delays associated with indemnity insurance.
Our main result is that, once these realistic frictions are introduced, parametric insurance can dominate indemnity insurance under the same utility objective. This dominance arises most clearly when households face binding budget constraints. In such cases, the deductible required to make indemnity insurance affordable may be so large that the contract is economically irrelevant. In contrast, a parametric contract can still provide a modest but meaningful transfer. We characterize indifference thresholds between the two designs and show that the welfare advantage of parametric insurance follows a non-monotonic pattern in the available premium budget: it emerges when budgets are small, disappears as indemnity insurance becomes effective, and vanishes entirely once both contracts are unconstrained.
@@ 4.3 Budget-constrained choice and two-dimensional budget surfaces
We now adopt the comparison most relevant for affordability: an agent faces an exogenous premium budget $\bar{P}$ and chooses the best contract they can buy within that budget. Critically, the agent can always opt out and remain uninsured. This formulation captures the practical asymmetry: under indemnity pricing with fixed costs, low budgets may translate into either infeasibility or an effectively useless deductible layer, whereas the same budget translates mechanically into a positive parametric payment as long as $\bar{P}>(1+\theta_{p})\gamma_{p}$.
@@
Figure 3(a) fixes $\gamma_{d}=1,\!000$ (and the remaining parameters) and plots, as functions of $\bar{P}$: (i) the effective premium actually spent by each product (which can be strictly below $\bar{P}$ once the unconstrained optimum becomes affordable), and (ii) the corresponding budget-constrained mean–variance values $MV^{(p)}_{\mathrm{bud}}(\bar{P})$ and $MV^{(d)}_{\mathrm{bud}}(\bar{P},\gamma_{d})$. For tiny budgets, indemnity may be infeasible due to the additive friction: the minimum feasible indemnity premium at deductible $L$ is $P_{d}^{\min}(\gamma_{d})=\mathcal{P}^{(d)}(L;\theta_{d},\gamma_{d})=(1+\theta_{d})\gamma_{d}.$ In that region, the best indemnity-or-no-insurance option is simply no insurance ($MV^{(0)}$), while parametric can already begin providing welfare gains if $\gamma_{p}$ is smaller and the linear premium constraint allows $k>0$.
@@ A.4 Mean–variance objective: deductible indemnity
$W=w_{0}-\mathcal{P}^{(\text{.})}\left(.;.\right)-S+\mathfrak{B}^{(\text{.})}\left(N,Y_{1},\dots,Y_{N},\cdot\right),\qquad\operatorname{MV}^{(\text{.})}\left(W;\beta\right)=\mathbb{E}[W]-\beta\,\mathrm{Var}(W),\quad\beta>0.$ (A.10)

### v4

@@ Abstract
In high-risk environments, traditional indemnity insurance is often unaffordable or ineffective, despite its well-known optimality under expected utility. We compare excess-of-loss indemnity insurance with parametric insurance within a common mean-variance framework, allowing for fixed costs, heterogeneous premium loadings, and binding budget constraints. We show that, once these realistic frictions are introduced, parametric insurance can yield higher welfare for risk-averse individuals, even under the same utility objective and without relying on behavioral assumptions. The welfare advantage arises precisely when indemnity insurance becomes impractical, and disappears once both contracts are unconstrained. Our results help reconcile classical insurance theory with the growing use of parametric risk transfer in high-risk settings.
@@ 1 Introduction
Insurance markets are increasingly strained in regions exposed to low-frequency, high-severity risks such as floods, wildfires, and cyclones. In many such areas, indemnity insurance has become prohibitively expensive or altogether unavailable, as rising hazard intensity, capital requirements, and fixed costs push premiums and deductibles to levels that leave households effectively uninsured. A key practical limitation is that household indemnity insurance is typically written on a full-value (sum-insured) basis: the premium is charged to cover the entire exposure, and the main lever to restore affordability is to raise the deductible. In high-risk areas, this mechanism can push deductibles to economically unrealistic levels (e.g., tens or hundreds of thousands of dollars), rendering the contract theoretically available but impractical, and delivering negligible effective protection. By contrast, parametric insurance is inherently scalable: households can purchase a smaller, explicitly bounded layer of protection (e.g., a fixed payment per event) without having to insure the full loss. Crucially, parametric coverage scales by adding first-dollar protection, whereas scaling indemnity insurance requires removing first-dollar coverage through higher deductibles, quickly rendering partial indemnity ineffective.
Recent developments in flood and wildfire insurance markets illustrate this (un)insurability trend: insurers have withdrawn from parts of the California residential market, flood insurance take-up remains persistently low in high-risk zones, and coverage gaps are expanding even in advanced economies (e.g., Kousky and Cooke, 2012; Jarzabkowski et al., 2023; UN Office for Disaster Risk Reduction, 2025). Yet when disasters occur, losses are rarely borne privately alone. Instead, they are often socialized ex post through government relief, reconstruction grants, and implicit guarantees. This growing disconnect between private insurance provision and public risk-bearing raises a fundamental economic question: when full indemnification is no longer feasible, what forms of risk transfer can still improve individual welfare?
Classical insurance theory provides a clear benchmark. Under broad and well-understood conditions, an excess-of-loss indemnity contract maximizes the expected utility of a risk-averse agent when premiums are loaded proportionally to expected losses (Arrow, 1971; Raviv, 1979). This result has shaped both theoretical and practical views of optimal insurance design for decades. However, it relies on assumptions that are increasingly violated in high-risk environments. In particular, it abstracts from differences in risk borne and fixed costs borne by insurers across contract types, as well as binding budget constraints faced by households. When insurance is expensive, the deductible that is still affordable becomes so large that indemnity insurance offers little or no effective protection, even though it remains optimal in a purely formal sense. In such cases, the classical optimality result continues to hold mathematically, but loses much of its economic content.
@@
Nevertheless, indemnity insurance is still widely considered the best (or only) way to provide effective protection. We show that this view is incomplete. Using a tractable model of frequency and severity, we compare indemnity and parametric insurance under the same expected mean–variance utility criterion. Losses arrive according to a Poisson process and have censored exponential severity. This specification can capture the salient features of natural disaster risk (low-frequency, high-severity) while allowing for closed-form solutions. The indemnity contract is of excess-of-loss type with a deductible, while the parametric contract pays a fixed amount per event and shares the same trigger, so there is no basis risk on frequency. Importantly, we allow for distinct premium loadings and fixed costs across contract types, reflecting the higher capital requirements, claims management costs, and reconstruction delays associated with indemnity insurance. At the sovereign level, parametric instruments are widely used as part of fiscal layering strategies, where rapid, rules-based payouts provide immediate liquidity after disasters while larger losses are addressed through slower mechanisms. A similar logic applies at the household level: even limited parametric payments can play a critical role in enabling early recovery and reducing reliance on ad hoc public assistance.
Our main result is that, once these realistic frictions are introduced, parametric insurance can dominate indemnity insurance under the same utility objective. This dominance arises most clearly when households face binding budget constraints. In such cases, the deductible required to make indemnity insurance affordable may be so large that the contract is economically irrelevant. In contrast, a parametric contract can still provide a modest but meaningful transfer. Even when this transfer covers only a small fraction of the realized loss, it can materially improve welfare by reducing liquidity shortfalls and supporting early recovery. We characterize indifference thresholds between the two designs and show that the welfare advantage of parametric insurance follows a non-monotonic pattern in the available premium budget: it emerges when budgets are small, disappears as indemnity insurance becomes effective, and vanishes entirely once both contracts are unconstrained.
@@ 3.2 Optimal results under censored exponential losses
3.3 Interpretation: tail coverage versus first-dollar coverage
The identity (B.5) admits a clear economic interpretation of how indemnity and parametric contracts allocate protection across the loss distribution. Under the optimal excess-of-loss indemnity contract, insurance coverage applies only to losses exceeding the deductible $d^{*}$: indemnity insurance therefore protects the tail of the loss distribution. By contrast, the parametric contract pays a fixed amount $k^{*}$ whenever an insured event occurs, independently of its severity. Parametric insurance thus provides first-dollar protection.
This distinction is largely irrelevant when contracts are unconstrained and risk is low. However, it becomes decisive once affordability or budget constraints are introduced. Scaling down a parametric contract by reducing $k$ proportionally reduces first-dollar protection while preserving its qualitative role: the insured continues to receive immediate liquidity whenever an event occurs. Scaling down an indemnity contract, by contrast, requires increasing the deductible $d$, which removes coverage precisely where losses are most likely to materialize. As a result, partial indemnity insurance rapidly becomes economically ineffective: very high deductibles leave households self-insuring most realizations, even though coverage remains formally available.
This asymmetry explains why budget constraints affect the two designs so differently. Parametric insurance remains scalable and meaningful at low budgets because it adds protection “from the ground up”. Indemnity insurance, on the other hand, can only become affordable by covering ever thinner slices of the tail of the loss distribution. In high-risk environments, where premiums are dominated by fixed costs and capital loadings, this mechanism renders partial indemnity impractical well before it becomes formally infeasible.
Finally, the provision of first-dollar protection is often viewed as undesirable in traditional insurance settings because of moral hazard concerns. In the context of disaster risk, however, this argument carries much less force. Households have no control over the occurrence of floods, storms, or wildfires, and therefore cannot influence the triggering of parametric payouts. Moreover, because the parametric payment is fixed, insured households retain strong—and arguably stronger—incentives to mitigate loss severity, as any reduction in damage directly benefits them. In this setting, the absence of a deductible is therefore not a drawback of parametric insurance, but a feature that enhances its effectiveness when only partial coverage is affordable.
@@ 4.3 Budget-constrained choice and two-dimensional budget surfaces
We now adopt the comparison most relevant for affordability: an agent faces an exogenous premium budget $\bar{P}$ and chooses the best contract they can buy within that budget. Critically, the agent can always opt out and remain uninsured. This formulation captures the practical asymmetry: under indemnity pricing with fixed costs, low budgets may translate into either infeasibility or an effectively useless deductible layer, whereas the same budget translates mechanically into a positive parametric payment as long as $\bar{P}>(1+\theta_{p})\gamma_{p}$. This pattern reflects the fact that parametric insurance provides first-dollar liquidity even at low budgets, whereas indemnity insurance becomes affordable only by pushing coverage into the extreme tail.
@@
Figure 3(a) fixes $\gamma_{d}=1,\!000$ (and the remaining parameters) and plots, as functions of $\bar{P}$: (i) the effective premium actually spent by each product (which can be strictly below $\bar{P}$ once the unconstrained optimum becomes affordable), and (ii) the corresponding budget-constrained mean–variance values $MV^{(p)}_{\mathrm{bud}}(\bar{P})$ and $MV^{(d)}_{\mathrm{bud}}(\bar{P},\gamma_{d})$. For tiny budgets, indemnity may be infeasible due to the additive friction: the minimum feasible indemnity premium at deductible $L$ is $P_{d}^{\min}(\gamma_{d})=\mathcal{P}^{(d)}(L;\theta_{d},\gamma_{d})=(1+\theta_{d})\gamma_{d}.$ In that region, the best indemnity-or-no-insurance option is simply no insurance ($MV^{(0)}$), while parametric can already begin providing welfare gains if $\gamma_{p}$ is smaller and the linear premium constraint allows $k>0$. Although parametric payouts may be small relative to extreme losses in this region, their welfare value arises precisely from providing immediate and certain resources when indemnity coverage is infeasible. The comparison is therefore second-best by construction, and the relevant benchmark is the no-insurance outcome rather than full loss replacement.
@@ 5.3 Speed of payout, recovery, and resilience
Furthermore, parametric insurance at the household level could also potentially mirror the logic of fiscal risk layering in sovereign disaster-risk management. Governments routinely rely on parametric or quasi-parametric instruments not to cover the full cost of disasters, but to secure rapid liquidity that stabilizes short-run fiscal and economic conditions. Rather than a replacement, parametric insurance could function as a liquidity layer that complements, rather than replaces, other available risk transfers.
@@ A.4 Mean–variance objective: deductible indemnity
$W=w_{0}-\mathcal{P}^{(\text{.})}\left(.;.\right)-S+\mathfrak{B}^{(\text{.})}\left(N,Y_{1},\dots,Y_{N},\cdot\right),$ (A.10)
and associated mean-variance
$\operatorname{MV}^{(\text{.})}\left(W;\beta\right)=\mathbb{E}[W]-\beta\,\mathrm{Var}(W),\quad\beta>0.$ (A.11)